The table below shows data on the country of origin and worries about climate change for respondents of a survey.

Country: Country A Country B Country C Country D

No worries 124 99 32 53

Some worries 12 44 76 235

Big worries 0 12 41 555

Compute Cramer’s V for the association between country of origin and worries about climate change. Show your calculations and interpret your result.
Answer #1

Answer:

Step-1 :

Chi-square calculation as below;

I. Observed counts
Country A Country B Country C Country D Column Total (CT)
No worries 124 99 32 53 308
Some worries 12 44 76 235 367
Big worries 0 12 41 555 608
Row Total (RT) 136 155 149 843 1283 Grand Total (GT)
II. Expected counts = ( RT x CT ) /GT
Expected frequency (E) Country A Country B Country C Country D Total
No worries                       32.65                       37.21                       35.77                         202.37 308
Some worries                       38.90                       44.34                       42.62                         241.14 367
Big worries                       64.45                       73.45                       70.61                         399.49 608
Total 136 155 149 843 1283
III. Chi-square components
Chi-square calculations Country A Country B Country C Country D Total
No worries                    255.60                    102.61                         0.40                         110.25                      468.86
Some worries                       18.60                         0.00                       26.14                              0.16                        44.90
Big worries                       64.45                       51.41                       12.42                            60.54                      188.82
Total                    338.66                    154.02                       38.95                         170.95       702.58 <== Chi-square value

Step-2:

Nominal Measures: Cramers V (V) Cramers V is used for tables larger than 2x2. Formula for Cramers V: x² V= V(N)(min r - 1,

Hence,

V=sqrt(702.58/(1283*2)) = 0.5233

Describing Strength of Association : Cramer's V

Characterizations

>.5                   high association

.3 to .5             moderate association

.1 to .3             low association

0 to .1              little if any association

Conclusion :

Since V = 0.5233 > 0.5 , it implies high association between country of origin and worries about climate change.

######## End of answer